OpenStudy (anonymous):

Consider the vector v = (0;3;4) in 3D. Why are there infinite vectors that are perpendicular to v in 3D?

OpenStudy (anonymous):

Because the vector is a line segment containing an infinite number of points - any of which may have a line (vector) drawn perpendicular to it.

OpenStudy (anonymous):

Should I answer this question by saying what you just said?

OpenStudy (anonymous):

I would answer the question the way I did. You should understand what my answer means and decide whether I'm correct or not rather than trust me to tell you whether you should quote me or not. I believe my concept is correct, but I don't know if I stated it in the terminology your teacher wants to hear.

OpenStudy (anonymous):

hi friends, can I say something?

OpenStudy (anonymous):

Sure. I don't rent the place out. :-)

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

you have the formula to consider whether 2 vectors are perpendicular or not. if they are, <u,v> =0

OpenStudy (anonymous):

sorry...just found that the actual question is why there are infinite UNIT vectors that are perpendicular to v in 3D

OpenStudy (anonymous):

your u = (0,3,4) let v =(a,b,c) <u,v> = 0a +3b+4c =0 you have 1 equation, 3 variables the solution for that is infinite set

OpenStudy (anonymous):

that is terminology of proving how to get and why it is infinite solution. does it make sense?

OpenStudy (anonymous):

but they are not necessarily unit vectors, right?

OpenStudy (anonymous):

i think so, it's in general, of course including unit vector.

OpenStudy (anonymous):

unit vector is just a vector with the length 1. that's it

OpenStudy (anonymous):

moreover, your vector u=(0,3,4) is not unit vector

OpenStudy (anonymous):

it's arbitrary vector, and no need to convert it to unit vector to consider

OpenStudy (anonymous):

yes i know...what i want to know is why there are infinite unit vectors perpendicular to v

OpenStudy (anonymous):

so, does mine make sense to you? 1 question, 3 variables, you have at least 2 free variables. base on that you can conclude the infinite solution for the stuff. I am sure.

OpenStudy (anonymous):

sounds quite right but if we have 0a +3b+4c =0, the equation can be reduced to 3b+4c =0, and so only 2 variables left

OpenStudy (anonymous):

and unit vector is of the length of 1, does it mean that we can set up another equation?

OpenStudy (anonymous):

hey friend, I think you misunderstanding something in that. cannot reduce because if you do that, you miss a. I mean with any a, 0a =0 . for example if a =1, 0a=0, a =2 ,0a =0, too, a =3; 0a =0 and so on,... that mean infinite set of a, not a =0 is only one solution. you know what I mean?

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

hey, how about the asker, where is he?

OpenStudy (anonymous):

that's me

OpenStudy (anonymous):

ok, sorry friend.

OpenStudy (anonymous):

as the unit vectors are of the length of 1, can we set up another equation ?

OpenStudy (anonymous):

particular problem, please

OpenStudy (anonymous):

as i said above, what i want to find out is UNIT vectors that are perpendicular to v=(0, 3, 4) in 3D can i let u=(a,b,c) and then set up the following equation? 1/(a^2+b^2+c^2)(sqrt)=1

OpenStudy (anonymous):

(a,b,c)/(a^2+b^2+c^2)(sqrt)=1

OpenStudy (anonymous):

in the original question, you just ask the proof of infinite set of vectors which perpendicular to (0,3,4). I give you that proof, think of this, all unit vectors =1, generalize the vector perpendicular (a,b,c) in the set you have \[\frac{ <a,b,c> }{ \sqrt{a^2+b^2+c^2} }\] is unit vector where <(0,3,4),(a,b,c)>=0 I think so

OpenStudy (anonymous):

sorry for my mistake...

OpenStudy (anonymous):

even if we have 2 equations now, there are still infinite unit vectors perpendicular to v, right?

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

@richard1024 exactly.. any combination of "b" and "c" that satisfies the above rule alongwith any real "a" will be the perpendicular to <u>

OpenStudy (anonymous):

@qweqwe123123123123111 said that in plain English, @hoa said the same exact thing in mathematical terms :) and I just restating them and connected them with a thread.

OpenStudy (anonymous):

we have these 2 equations (a,b,c)/(a^2+b^2+c^2)(sqrt)=1 and 0a+3b+4c=0

OpenStudy (anonymous):

still there can be infinite solutions?

OpenStudy (anonymous):

they both give the same equation. you wrote the first one incorrectly.

OpenStudy (anonymous):

the numerator should be <(0,3,4),(a,b,c)>

OpenStudy (anonymous):

which is the same exact thing as the other equation. so, you've got only one equation. kapeesh?

OpenStudy (anonymous):

why they both give the same equation?

OpenStudy (anonymous):

eq# 1 <(0,3,4),(a,b,c)>=0 eq# 2 <(0,3,4),(a,b,c)> --------------- = 0 sqrt(a^2+b^2+c^2)

OpenStudy (anonymous):

now, aint they the same thing?

OpenStudy (anonymous):

eq. 2 is just the unit vectors..

OpenStudy (anonymous):

of the equation 1

OpenStudy (anonymous):

i have just attached a file are you able to open it?

OpenStudy (anonymous):

what i don't know is question 1b(iii)

OpenStudy (anonymous):

yes. that is infinite.

OpenStudy (anonymous):

i don't understand why we have this equation <(0,3,4),(a,b,c)> --------------- = 0 sqrt(a^2+b^2+c^2)

OpenStudy (anonymous):

then just forget it.. its misleading.. wipe it off... :)

OpenStudy (anonymous):

its not needed

OpenStudy (anonymous):

just remember <(0,3,4),(a,b,c)>

OpenStudy (anonymous):

as it is a unit vector, should we have an equation like this instead? (a,b,c) --------------- = 1 sqrt(a^2+b^2+c^2)

OpenStudy (anonymous):

why it's (0,3,4),(a,b,c)?

OpenStudy (anonymous):

well, thats the last step.. you are stuck in an infinite loop there... you first have to snap out of it

OpenStudy (anonymous):

now i understand why (a,b,c) --------------- . (0,3,4)= 0 sqrt(a^2+b^2+c^2)

OpenStudy (anonymous):

and unit vector has a length of 1, so is the following equation correct? (a,b,c) --------------- = 1 sqrt(a^2+b^2+c^2)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[(a,b,c)\div \sqrt{a ^{2}+b ^{2}+c ^{2}}=1\]

OpenStudy (anonymous):

nice work

OpenStudy (anonymous):

it's correct?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

that's why i said we have 2 equations

OpenStudy (anonymous):

ok. but they are "parallel" and hence, same but you got the idea.

OpenStudy (anonymous):

but still there are infinite solutions because we have 2 equations but 3 unknowns, right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

now i understand everything thank you very much for your patience, time, and all you told me

OpenStudy (anonymous):

you are welcome. in math, when you get stuck somewhere, snap out, clear your mind and look at it with fresh eyes.

OpenStudy (anonymous):

okay :)